OpenStudy (babtaooche):

Help, I'll fan and medal Find f(x) and g(x) so that the function can be described as y = f(g(x)). y = two divided by x squared + 9

OpenStudy (michele_laino):

here is an example: \[\begin{gathered} g\left( x \right) = {x^2} \hfill \\ f\left( x \right) = \frac{2}{{x + 9}} \hfill \\ \end{gathered} \]

OpenStudy (michele_laino):

as we can see, we have: \[\left( {g\left( x \right)} \right) = \frac{2}{{g\left( x \right) + 9}} = ...?\] please complete

OpenStudy (michele_laino):

oops... \[f\left( {g\left( x \right)} \right) = \frac{2}{{g\left( x \right) + 9}} = ...?\]

OpenStudy (babtaooche):

I tried but I didn't get an answer..

OpenStudy (michele_laino):

try to substitute \(g(x)=x^2\), into my last formula above

OpenStudy (babtaooche):

Ugh, I just got 0

OpenStudy (michele_laino):

hint: if I replace \(g(x)\) with its value \(x^2\), I get: